i need to solve this equation for d.
4(d-3)=2d

Answer by ankor@dixie-net.com(22740)

solve this equation for d.
4(d-3) = 2d
perform a few simple algebra operations
4d - 12 = 2d
4d = 2d + 12
4d - 2d = 12
2d = 12
d = 12/2
d = 6
:
:
Check in the original equation, replacing d with 6
4(6-3) = 2(6)
4(3) = 12